OpenStudy (anonymous):

Use square too property: 3z^2-63=0 the 63...err i got lost there :( Do i square it or what?

OpenStudy (anonymous):

3z^2 = 63 z^2 = 21 z = \[\pm \sqrt{21}\]

OpenStudy (anonymous):

how did you come by the 21? thanks :)

OpenStudy (mathteacher1729):

Moneybird took 3z^2 = 63 divide both sides by 3 z^2 = 21 Take square roots of both sides z = sqrt(21) and z = -sqrt(21)